# Negation-Aware Test-Time Adaptation for Vision-Language Models

### v1

@@ Abstract
In this paper, we study a practical but less-touched problem in Vision-Language Models (VLMs), i.e., negation understanding. Specifically, many real-world applications require models to explicitly identify what is false or non-existent, e.g., radiologists may search for images that exclude specific conditions. Despite the impressive transferability of VLMs through large-scale training, they suffer from a critical limitation that fails to handle negation. To address this challenge, existing methods attribute its root cause to the scarcity of negation training data and propose to fine-tune VLMs on massive data containing explicit negation. Undoubtedly, such data-centric solutions demand substantial data and computational resources, limiting their sustainable widespread adoption. To tackle negation in a low-carbon manner, we empirically observe that the key obstacle lies in the dual-concept shifts between the affirmation and negation distributions. Therefore, we propose a Negation-Aware Test-Time Adaptation (NEAT) method to efficiently adjust distribution-related parameters during inference. In brief, NEAT can reduce distribution shift in consistent semantics while eliminating false distributional consistency in unrelated semantics. Extensive experiments on the various negation understanding tasks verify the effectiveness of the proposed method. The code is available at https://github.com/hhc1997/NEAT.
@@ I Introduction
Figure 1: A toy example to illustrate the dual-concept shifts problem in Vision-Language Models when understanding negation. CLIP learns diverse open-world knowledge but struggles with negation contexts. And we show how distributional adaptation can improve its performance.
Based on the above discussions and observations, instead of the full-parameter fitting on negation data, we propose only adjusting distribution-related parameters (e.g., normalization layers) to tackle the dual-concept shifts during inference. To be specific, we propose a novel Negation-Aware Test-time Adaptation (dubbed NEAT) framework, which can rapidly learn from unlabeled multi-modal negation data and then make reliable predictions. Our NEAT encapsulates three key components. First, the negation-separated refinement module is adopted to obtain refined entropy between negated captions and visual content, where descriptions are decomposed into two affirmative parts for VLMs to handle adeptly. Second, to address the false distributional consistency in unrelated semantics, we propose to view the reversed negation-conditioned caption as the hardest negative information that shares opposite semantics to its corresponding visual sample. Theoretically, we show that this reversed contrastive object only requires comparing against limited visual samples and is equivalent to a simple metric loss. Third, to achieve efficient adaptation, we suggest directly reducing the textual-distribution gap arising from dual-concept shifts. Empirically, NEAT establishes state-of-the-art performance on multiple negation benchmarks that span images, videos, and medical scenarios, outperforming the expensive post-trained baselines by a significant margin and observing strong generalization ability in adapted VLMs.
@@
  • To the best of our knowledge, our paper pioneers the exploration of test-time adaptation to enhance negation understanding of VLMs. We reveal that the key to negation understanding is the dual-concept shifts problem between affirmation and negation distributions, which could be addressed by adapting lightweight normalization layers of models.
@@ II-A Negation Understanding in Vision-Language Models
Vision-language foundation models learned diverse knowledge from large-scale multi-modal data, which have attracted significant attention due to their powerful transfer capabilities. By supervising visual representations with natural language descriptions, pioneering works CLIP [31] and ALIGN [17] show great success in various downstream tasks. However, the web-crawled training data are mainly affirmative, which makes it difficult to comprehend the negation of the foundation VLMs. To comprehensively evaluate model performance in negation scenarios, recent works CREPE [25] and CC-Neg [33] proposed vision-language benchmarks that target compositional understanding with negation, but the reliance on linguistic templates limits the diversity of real negation queries. In comparison, NegBench [2] utilizes LLMs to generate more natural negated captions that span images, videos, and medical datasets. Despite being trained on billion-scale data, the unsatisfactory performance on these benchmarks reveals that VLMs struggle to understand negation.
@@ II-B Test-Time Adaptation
The distribution shift between training and test data poses a key challenge when transferring the zero-shot capabilities of foundation VLMs. In practice, such shifts are inevitable due to environmental variations [19] or the encounter of unseen concepts [38]. To address the problem, considerable efforts have been devoted to developing training-time solutions: domain adaptation methods [12] that narrow gaps between source and target samples during training. Domain generalization methods [1] that directly learn domain-invariant representations through robust training strategies. However, these methods may require access to source domain data and cannot achieve adaptation in an online manner, limiting their practicality in real-time applications such as autonomous driving.
@@
Figure 2: Dual-concept shifts problem observed in pre-trained VLMs when understanding negation. Distributions of normal text, negation-conditioned text, and reversed negation-conditioned text of the test set of CIFAR10 are shown on the left, middle, and right, respectively. All embeddings are produced by the OpenAI CLIP and visualized with UMAP [26] for dimension reduction. The Maximum Mean Discrepancy (MMD) metrics are also provided above for quantitative analysis. The distributions of normal and negation-conditioned text show a strong shift (MMD=1.571.571.57) despite sharing the same semantics. Meanwhile, the distributions of negation-conditioned and its reversed text show high consistency (MMD=0.080.080.08) despite being semantically opposite.
@@ IV-B1 Entropy Refinement via Negation Separation
where $\alpha$ is a trade-off parameter to control the penalty and $[x]_{+}=max(x,0)$ is the hinge function. Intuitively, higher $S(\hat{\bm{t}}_{i},\bm{v}_{j})$ requires ${v}_{j}$ depicts $\hat{t}_{i}^{P}$ well and remains unrelated to concepts in $\hat{t}_{i}^{N}$. Then we select the most similar sample by Eq.(4) to obtain the corresponding visual candidates $\bm{V}_{test}^{R}$. Consequently, the entropy minimizing objective for the online batch is
@@ IV-B2 Semantics Reversion Learning
where the pair $(\bar{t},\hat{v})$ is sampled from the joint distribution $p(\bar{t},\hat{v})$ and the pair $(\bar{t},\hat{v}^{\prime})$ is built by samples independently sampled from the marginal $p(\hat{v}^{\prime})$, and $\tau_{2}$ is a temperature parameter.
@@ V-A2 Implementation Details
TABLE III: Comparisons with state-of-the-art only TTA methods on MS-COCO and MSR-VTT datasets under negated queries. All methods are based on the ViT-B/32 CLIP. The best results are marked in bold.
@@
TABLE IV: Comparisons with state-of-the-art only TTA methods on MS-COCO and MSR-VTT datasets under negated queries. All methods are based on the ViT-B/32 NegCLIP. The best results are marked in bold.
@@ V-B1 Compared to Post-training Methods
  • Our NEAT could significantly improve the negation comprehension ability of pre-trained VLMs, e.g., it improves the rSum of CLIP by 11.80$\sim$29.54on MS-COCO.
  • Our NEAT could generalize to different VLMs and modalities. For instance, it improves the rSum of NegCLIP and its CC12M fine-tuned model by 10.8% and 10.2% on MS-COCO, respectively. In addition, although only adapted on 1,000 video-text pairs, NEAT still improves the rSum of CLIP by 7.0$\sim$16.5on MSR-VTT.
@@
TABLE V: Zero-shot transfer evaluation on two template-based negation classification tasks. The best results are marked in bold.
@@ V-B2 Compared to Test-time Adaptation Methods
[…]
@@
TABLE VI: Zero-shot transfer evaluation of TTA methods on the VOC2007 MCQ-Neg task after adaptation on unlabeled MS-COCO Retrieval-Neg data.
@@
V-C Generalizability Analysis
V-C1 Easy: Cross-Dataset Generalization
We first evaluate VLMs that were adapted by different TTA methods on the unseen VOC2007 MCQ-Neg task. As demonstrated in Table VI, although some TTA baselines can adapt to negation data and enhance performance on such data, the adapted models fail to generalize to novel negation data, leading to substantial performance drops. In contrast, our NEAT remarkably improves CLIP from 38.70% to 55.54% and NegCLIP from 30.47% to 49.73% on the unseen voc2007 dataset.
V-C2 Normal: Cross-Task Generalization
We further test NEAT-adapted VLMs across 9 widely used image classification datasets using two template-based negation understanding tasks. As elaborated in IV-A, we randomly introduce distractor labels to construct the negation-conditioned caption, and swap its two class names to construct the reversed negation-conditioned caption. The comparison results are shown in Table V. From the table, one could see that most negation-aware methods cannot generalize to downstream image classification tasks, performing even worse than the base model. Additionally, post-training on millions of well-designed negation data can endow VLMs with strong negation generalization capabilities, but the performance remains sub-optimal for the more difficult reversed negation-conditioned task. Remarkably, although only updating LN layers with minimal unlabeled data, our approach not only improves the accuracy of CLIP from 54.57% to 63.18% (average) for the negation-conditioned task, but also reduces the error rate of CLIP from 11.42% to 1.36% (average) for the reversed negation-conditioned task.
(a) Negation-separated Prediction
(b) Parameter Ablation
(c) Temperature Sensitivity
TABLE VII: Zero-shot transfer evaluation on CheXpert by replacing LN layers from different negation-aware models.
@@ V-C3 Hard: Cross-Normalization Layers Generalization
Beyond validating on complete models, we further investigate whether the adapted LN layers can be directly applied to domain-specific models. To this end, we use the medical foundation VLM, i.e., PLIP [15], as the base model, replacing its textual LN layers with those from negation-aware VLMs to observe the performance impact. Following [2], we conduct evaluations on two classification settings. Specifically, the Binary Classification-Aff task requires the model to correctly associate each image with the two affirmative statements, such as ‘This image shows Atelectasis’ and ‘This image shows Consolidation’. The Binary Classification-Neg task requires the model to distinguish each image with the presence or absence of a medical condition, such as ‘This image shows Atelectasis’ and ‘This image does not show Atelectasis’. As shown in Table VII, one can observe that PLIP exhibits robust negation understanding abilities in most disease recognition tasks, e.g., Atelectasis and Consolidation, suggesting that its training medical reports likely contain explicit negation expressions. Second, using the LN layers from post-training methods does not lead to significant performance changes, indicating that their negation understanding capabilities are more attributable to the encoder or decoder layers. Third, using the LN layers adapted by TCR brings performance gains in certain scenarios, but exhibits instability in others. Notably, our NEAT provides stable performance improvements overall, especially for the most frequent [2] condition Lung Opacity, i.e., achieving 14.54% and 36.11% improvements for Binary Classification-Aff and Classification-Neg tasks, respectively.
TABLE VIII: Ablation study of the proposed training objects.
$\mathcal{L}_{ent}$ $\mathcal{L}_{sr}$ $\mathcal{L}_{tri}$ I2T R@1 T2I R@1 MCQ
45.30 24.96 39.25
48.06 29.20 28.69
7.12 4.67 42.78
49.74 28.95 27.26
44.96 30.05 45.91
48.98 30.00 48.41

### v2

@@ Abstract
In this paper, we study a practical but less-touched problem in Vision-Language Models (VLMs), i.e., negation understanding. Specifically, many real-world applications require models to explicitly identify what is false or non-existent, e.g., radiologists may search for images that exclude specific conditions. Despite the impressive transferability of VLMs through large-scale training, they suffer from a critical limitation that fails to handle negation. To address this challenge, existing methods attribute its root cause to the scarcity of negation training data and propose to fine-tune VLMs on massive data containing explicit negation. Undoubtedly, such data-centric solutions demand substantial data and computational resources, limiting their sustainable widespread adoption. To tackle negation in a low-carbon manner, we empirically observe that the key obstacle lies in the dual-concept shifts between the affirmation and negation distributions. Therefore, we propose a Negation-Aware Test-Time Adaptation (NEAT) method to efficiently adjust distribution-related parameters during inference. In brief, NEAT can reduce distribution shift in consistent semantics while eliminating false distributional consistency in unrelated semantics. Extensive experiments on the various negation understanding tasks verify the effectiveness of the proposed method. Remarkably, with less than 0.01% of trainable parameters, NEAT achieves comparable or superior performance to state-of-the-art post-training approaches. Our code is available at https://github.com/hhc1997/NEAT.
@@ I Introduction
Figure 1: A toy example to illustrate the dual-concept shifts problem in Vision-Language Models when understanding negation. Negation-conditioned (NC) texts negate absent elements and share consistent semantics with affirmative ones. Reversed negation-conditioned (RNC) texts wrongly affirm absent elements and negate present ones, which are semantically opposite to NC texts. Although CLIP learns diverse open-world knowledge, it exhibits a similarity gap between affirmative and NC texts, and high similarity between NC and RNC texts. We show how distributional adaptation can tackle this issue.
Based on the above discussions and observations, instead of the full-parameter fitting on negation data, we propose only adjusting distribution-related parameters (e.g., normalization layers) to tackle the dual-concept shifts during inference. To be specific, we propose a novel Negation-Aware Test-time Adaptation (dubbed NEAT) framework, which can rapidly learn from unlabeled multi-modal negation data and then make reliable predictions. Our NEAT encapsulates three key components. First, the negation-separated refinement module is adopted to obtain refined entropy between negated captions and visual content, where descriptions are decomposed into two affirmative parts for VLMs to handle adeptly. Second, to address the false distributional consistency in unrelated semantics, we propose to view the reversed negation-conditioned caption as the hardest negative information that shares opposite semantics to its corresponding visual sample. Theoretically, we show that this reversed contrastive objective only requires comparing against limited visual samples and is equivalent to a simple metric loss. Third, to achieve efficient adaptation, we suggest directly reducing the textual-distribution gap arising from dual-concept shifts. Empirically, NEAT establishes state-of-the-art performance on multiple negation benchmarks that span images, videos, and medical scenarios, outperforming the expensive post-trained baselines by a significant margin and observing strong generalization ability in adapted VLMs.
@@
  • To the best of our knowledge, this work could be the first study to enhance negation understanding of VLMs via test-time adaptation. We reveal that the key to negation understanding is the dual-concept shifts problem between affirmation and negation distributions, which could be addressed by adapting lightweight normalization layers of models.
@@ II-A Negation Understanding in Vision-Language Models
Vision-language foundation models learned diverse knowledge from large-scale multi-modal data, which have attracted significant attention due to their powerful transfer capabilities [49, 3]. By supervising visual representations with natural language descriptions, pioneering works CLIP [35] and ALIGN [19] show great success in various downstream tasks. However, the web-crawled training data are mainly affirmative, which makes it difficult to comprehend the negation of the foundation VLMs. To comprehensively evaluate model performance in negation scenarios, recent works CREPE [29] and CC-Neg [37] proposed vision-language benchmarks that target compositional understanding with negation, but the reliance on linguistic templates limits the diversity of real negation queries. In comparison, NegBench [2] utilizes LLMs to generate more natural negated captions that span images, videos, and medical datasets. Despite being trained on billion-scale data, the unsatisfactory performance on these benchmarks reveals that VLMs struggle to understand negation. Recent work [51] shows that such problems persist even in powerful MLLMs like GPT-4o and Claude 3.5, which suffer from hallucinations when processing user-provided negation arguments.
@@ II-B Test-Time Adaptation
The distribution shift between training and test data poses a key challenge when transferring the zero-shot capabilities of foundation VLMs [39, 13, 28]. In practice, such shifts are inevitable due to environmental variations [21] or the encounter of unseen concepts [43]. To address the problem, considerable efforts have been devoted to developing training-time solutions: domain adaptation methods [14] that narrow gaps between source and target samples during training. Domain generalization methods [1] that directly learn domain-invariant representations through robust training strategies. However, these methods may require access to source domain data and cannot achieve adaptation in an online manner, limiting their practicality in real-time applications such as autonomous driving.
@@
Figure 2: Dual-concept shifts problem observed in pre-trained VLMs when understanding negation. Distributions of normal text, negation-conditioned text, and reversed negation-conditioned text of the test set of CIFAR10 are shown on the left, middle, and right, respectively. All embeddings are produced by the OpenAI CLIP and visualized with UMAP [30] for dimension reduction. The Maximum Mean Discrepancy (MMD) metrics are also provided above for quantitative analysis. The distributions of normal and negation-conditioned text show a strong shift (MMD=$1.57$) despite sharing the same semantics. Meanwhile, the distributions of negation-conditioned and its reversed text show high consistency (MMD=$0.08$) despite being semantically opposite.
@@ IV-B1 Entropy Refinement via Negation Separation
where $\alpha$ is a trade-off parameter to control the penalty and $x]_{+}=max(x,0)$ is the hinge function. Intuitively, higher $S(\hat{\bm{t}}_{i},\bm{v}_{j})$ requires ${v}_{j}$ depicts $\hat{t}_{i}^{P}$ well and remains unrelated to concepts in $\hat{t}_{i}^{N}$. Then we select the most similar sample by Eq.(4) to obtain the corresponding visual candidates $\bm{V}_{test}^{R}$. Consequently, the entropy minimizing objective for the online batch is
@@ IV-B2 Semantics Reversion Learning
[…]
@@ V-A2 Implementation Details
TABLE III: Comparisons with state-of-the-art online TTA methods on MS-COCO and MSR-VTT datasets under negated queries. All methods are based on the ViT-B/32 CLIP. The best results are marked in bold.
@@
TABLE IV: Comparisons with state-of-the-art online TTA methods on MS-COCO and MSR-VTT datasets under negated queries. All methods are based on the ViT-B/32 NegCLIP. The best results are marked in bold.
@@
TABLE V: Comparisons with state-of-the-art offline TTA methods on MS-COCO datasets under negated queries. All methods are based on the ViT-B/16 BLIP. The best results are marked in bold.
Method Retrieval Neg Task Zero-shot MCQ-Neg Task
Image-to-Text Text-to-Image rSum MCQ Type Total
R@1 R@5 R@10 R@1 R@5 R@10 Affirmed Negated Hybrid
Base model: ViT-B/16 BLIP
BLIP 52.60 77.90 85.76 33.58 58.96 69.68 378.48 33.81 16.68 5.12 18.63
TENT 37.96 66.98 79.50 40.54 67.16 78.33 $370.47{{\downarrow_{8.01}}}$ 19.14 20.70 2.68 $14.03{{\downarrow_{4.60}}}$
SAR 37.24 67.74 78.94 40.35 68.06 78.04 $370.37{{\downarrow_{8.11}}}$ 11.71 23.53 1.09 $11.84{{\downarrow_{6.79}}}$
READ 55.24 79.56 87.36 36.32 62.90 73.27 $394.65{{\uparrow_{16.17}}}$ 35.73 16.52 5.62 $19.41{{\uparrow_{0.78}}}$
COME 44.84 70.80 79.64 37.11 63.78 74.17 $370.34{{\downarrow_{8.14}}}$ 31.74 17.59 3.68 $17.72{{\downarrow_{0.91}}}$
TCR 48.38 76.10 85.34 46.10 73.26 82.39 $411.57{{\uparrow_{33.09}}}$ 20.62 31.76 5.47 $18.99{{\uparrow_{0.36}}}$
*NEAT 59.22 83.00 89.98 41.12 68.40 78.47 $\textbf{420.19}{{\uparrow_{41.71}}}$ 77.21 28.34 49.95 $\textbf{52.49}{{\uparrow_{33.86}}}$
Base model: ViT-B/16 BLIP Fine-tuned on MS-COCO Training Data
BLIP 69.98 90.98 95.00 54.69 79.86 87.34 477.85 50.69 12.94 16.65 27.17
TENT 68.14 89.46 94.10 57.44 82.02 88.85 $480.01{{\uparrow_{2.16}}}$ 47.79 19.36 28.28 $32.16{{\uparrow_{4.99}}}$
SAR 66.40 88.10 93.50 57.51 82.13 88.90 $476.54{{\downarrow_{1.31}}}$ 47.29 19.36 29.97 $32.57{{\uparrow_{5.40}}}$
READ 70.40 91.28 95.22 53.01 78.62 86.34 $474.87{{\downarrow_{2.98}}}$ 53.00 11.76 17.20 $27.78{{\uparrow_{0.61}}}$
COME 69.38 90.62 94.96 57.63 81.81 88.75 $483.15{{\uparrow_{5.30}}}$ 55.81 11.12 16.90 $28.44{{\uparrow_{1.27}}}$
TCR 68.58 89.66 94.12 58.78 83.10 89.66 $483.90{{\uparrow_{6.05}}}$ 47.98 19.52 30.07 $32.89{{\uparrow_{5.72}}}$
*NEAT 74.18 92.40 95.98 56.40 81.11 88.47 $\textbf{488.54}{{\uparrow_{10.69}}}$ 58.27 18.29 46.22 $\textbf{41.53}{{\uparrow_{14.36}}}$
@@ V-B1 Compared to Post-training Methods
[…]
  • The normalization layers modified by NEAT do not impair VLMs’ ability for normal vision-language comprehension on affirmative statements, with rSum performance remaining within stable and acceptable ranges, i.e. $4.75\%\sim+4.91\%$.
@@
TABLE VI: Zero-shot transfer evaluation on two template-based negation classification tasks. The best results are marked in bold.
@@ V-B2 Compared to Test-time Adaptation Methods
  • Although some multi-modal methods show considerable performance gains on Retrieval-Neg Task, i.e., TCR, the accuracy on the corresponding MCQ-Neg task drops unexpectedly by $1.10\%$ to $11.54\%$. This indicates that the observed improvements may be attributed to reducing modality gaps rather than addressing the concept shift problem arising from negation understanding.
@@
In addition to the evaluation for CLIP-based VLMs, we also conduct experiments for BLIP-based VLMs [26], i.e., a base model pretrained on 129 million data and its fine-tuned model on the MS-COCO dataset. From Fig. V, one could see that some TTA baselines show large performance variations between image-to-text and text-to-image retrieval after adaptation. This may be because BLIP uses multiple pretraining objectives while adaptation only focuses on the image-text contrastive one. The phenomenon is significantly alleviated when the model is fine-tuned on corresponding image-text retrieval tasks. Moreover, like the observations on CLIP-based models, our NEAT remarkably boosts the effectiveness of the baselines. Specifically, NEAT improves BLIP by 41.71% (rSum) for cross-modal retrieval and 33.86% (accuracy) for multiple choice questions.
V-C Generalizability Analysis
V-C1 Easy: Cross-Dataset Generalization
We first evaluate VLMs that were adapted by different TTA methods on the unseen VOC2007 MCQ-Neg task. As demonstrated in Table VII, although some TTA baselines can adapt to negation data and enhance performance on such data, the adapted models fail to generalize to novel negation data, leading to substantial performance drops. In contrast, our NEAT remarkably improves CLIP from 38.70% to 55.54% and NegCLIP from 30.47% to 49.73% on the unseen voc2007 dataset.
V-C2 Normal: Cross-Task Generalization
We further test NEAT-adapted VLMs across 9 widely used image classification datasets using two template-based negation understanding tasks. As elaborated in IV-A, we randomly introduce distractor labels to construct the negation-conditioned caption, and swap its two class names to construct the reversed negation-conditioned caption. The comparison results are shown in Table VI. From the table, one could see that most negation-aware methods cannot generalize to downstream image classification tasks, performing even worse than the base model. Additionally, post-training on millions of well-designed negation data can endow VLMs with strong negation generalization capabilities, but the performance remains sub-optimal for the more difficult reversed negation-conditioned task. Remarkably, although only updating LN layers with minimal unlabeled data, our approach not only improves the accuracy of CLIP from 54.57% to 63.18% (average) for the negation-conditioned task, but also reduces the error rate of CLIP from 11.42% to 1.36% (average) for the reversed negation-conditioned task.
(a) Negation-separated Prediction
(b) Parameter Ablation
(c) Temperature Sensitivity
TABLE VII: Zero-shot transfer evaluation of TTA methods on the VOC2007 MCQ-Neg task after adaptation on unlabeled MS-COCO Retrieval-Neg data.
@@ V-C2 Normal: Cross-Task Generalization
TABLE VIII: Zero-shot transfer evaluation on CheXpert by replacing LN layers from different negation-aware models.
@@ V-C3 Hard: Cross-Normalization Layers Generalization
Beyond validating on complete models, we further investigate whether the adapted LN layers can be directly applied to domain-specific models. To this end, we use the medical foundation VLM, i.e., PLIP [17], as the base model, replacing its textual LN layers with those from negation-aware VLMs to observe the performance impact. Following [2], we conduct evaluations on two classification settings. Specifically, the Binary Classification-Aff task requires the model to correctly associate each image with the two affirmative statements, such as ‘This image shows Atelectasis’ and ‘This image shows Consolidation’. The Binary Classification-Neg task requires the model to distinguish each image with the presence or absence of a medical condition, such as ‘This image shows Atelectasis’ and ‘This image does not show Atelectasis’. As shown in Table VIII, one can observe that PLIP exhibits robust negation understanding abilities in most disease recognition tasks, e.g., Atelectasis and Consolidation, suggesting that its training medical reports likely contain explicit negation expressions. Second, using the LN layers from post-training methods does not lead to significant performance changes, indicating that their negation understanding capabilities are more attributable to the encoder or decoder layers. Third, using the LN layers adapted by TCR brings performance gains in certain scenarios, but exhibits instability in others. Notably, our NEAT provides stable performance improvements overall, especially for the most frequent [2] condition Lung Opacity, i.e., achieving 14.54% and 36.11% improvements for Binary Classification-Aff and Classification-Neg tasks, respectively.